# Vojta’s abc conjecture for entire curves in toric varieties highly ramified over the boundary

### v1

@@ 2.1. The Rational Case
(2.7) $N(G_{0}(u_{1},\ldots,u_{n}))-N^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\leq 2\epsilon\deg\mathbf{f}.$
@@ 2.2. Motivation for the proof of the Main Theorem
As in Section 2.1, the starting point of the proof of Theorem 6 is as the follows: For a given rational map, we define a set $S$ (depending on the map) that contains the zeros and poles of each coordinate function of the map. In this way, all the coordinate functions become $S$-units, and so we can apply Theorem 5. Importantly, the constants and the exceptional set obtained in Theorem 5 do not depend on $S$, allowing this method to be used for all rational maps. To apply this view point to the analytic case, we consider $f=(f_{0},\ldots,f_{n})$, where $f_{i}$, $0\leq i\leq n$, are entire functions without common zeros. Let $\mathcal{E}_{f}$ denote the set of all the zeros of $f_{i}$, $0\leq i\leq n$, and define $\mathcal{Y}_{f}:=\mathbb{C}\setminus\mathcal{E}_{f}$. Then ${\mathcal{Y}_{f}}$ is a parabolic open Riemann surface. On ${\mathcal{Y}_{f}}$, the restriction $\tilde{f}_{i}:=f_{i}|_{\mathcal{Y}_{f}}$ becomes a unit on ${\mathcal{Y}}$ (i.e. it has neither zeros nor poles). This perspective was also adopted in [2], where they laid important foundational work. For our purposes, we will need to develop an analogue of Theorem 5 over parabolic open Riemann surfaces.
@@ 3. Nevanlinna theory and the GCD theorem on parabolic Riemann surfaces
3.1. Parabolic Riemann surfaces
@@
1. For a hypersurface $D=\{Q=0\}$ in $\mathbb{P}^{n}({\mathbb{C}})$, where $Q$ is a homogeneous polynomial of degree $d$, the proximity function is defined as
[…]
$\lambda_{D}({\bf x})=\log{\|{\bf x}\|^{d}\over|Q({\bf x})|},~{}~{}{\bf x}\in\mathbb{P}^{n}({\mathbb{C}}).$
2. The characteristic (order) function is defined by
$T_{f,{\mathcal{Y}}}(r)=\int_{1}^{r}{dt\over t}\int_{B(t)}f^{*}\omega_{FS}.$
3. The counting function and the $k$-truncated counting function are defined by
$N_{f,{\mathcal{Y}}}(D,r):=\int_{1}^{r}\sum_{z\in B(t)}{\rm ord}_{z}f^{*}D{dt\over t},~{}~{}~{}~{}~{}~{}~{}\mbox{and}~{}~{}N_{f,{\mathcal{Y}}}^{(k)}(D,r):=\int_{1}^{r}\sum_{z\in B(t)}\mathop{\mathrm{min}}\{k,{\rm ord}_{z}f^{*}D\}{dt\over t}.$
4. The counting function extends naturally to the sub-schemes, as well as the notion $N_{\text{gcd},{\mathcal{Y}}}(f,g)(r)$, where $f,g$ are meromorphic functions on $\mathcal{Y}$. See [14].
Parabolic First Main Theorem. Let $D$ be a hypersurface in ${\mathbb{P}}^{n}({\mathbb{C}})$ and $f:{\mathcal{Y}}\rightarrow\mathbb{P}^{n}({\mathbb{C}})$ be a holomorphic map with $f({\mathcal{Y}})\not\subset D$. Then
$T_{f,{\mathcal{Y}}}(r)=m_{f,{\mathcal{Y}}}(D,r)+N_{f,{\mathcal{Y}}}(D,r)+O(1).$
It is important to note that, when we write $O(g(x))$ for some function $g$, we mean that $Cg(x)$ where $C$ is a constant independent of ${\mathcal{Y}}$ throughout the paper.
Fix a nowhere vanishing global holomorphic vector field $\xi\in\Gamma({\mathcal{Y}},T_{\mathcal{Y}})$ (we denote it by ${\partial\over\partial z}$) on ${\mathcal{Y}}$. Such a vector field exists because ${\mathcal{Y}}$ is non-compact and consequently its holomorphic tangent bundle $T_{\mathcal{Y}}$ is holomorphically trivial. From Lemma 4.4 in [10],
$\int_{S_{r}}\log\left|d\sigma\left({\partial\over\partial z}\right)\right|^{2}d\mu_{r}=-\int_{1}^{r}\chi_{\sigma}(t)\frac{dt}{t}+2\varsigma\log r+O(1),$
where $\chi_{\sigma}(r)$ is the Euler characteristic of $B(r)$, and
[…]
which is independent of $r$ for $r$ being sufficiently large. Denote by
$\mathfrak{X}^{+}_{\sigma}(r):=\int_{S(r)}\log^{+}\left|d\sigma\left({\partial\over\partial z}\right)\right|^{2}d\mu_{r}.$
[…]
(3.1) $\hat{\tau}:=\log\hat{\sigma}:=\log^{+}|z|+\sum_{j\geq 1}r_{j}\log^{+}{r_{j}\over|z-a_{j}|}.$
Note that $\hat{\tau}$ takes values in $[0,+\infty)$, is continuous on ${\mathcal{Y}}$ and is smooth (indeed harmonic) outside the circle $S(0,1):=\{|z|=1\}$ and the disjoint circles $S(a_{j},r_{j}):=\{|z-a_{j}|=r_{j}\}$. By the Poincare-Lelong formula, it is clear that
$dd^{c}\hat{\tau}={1\over 2}\nu(0,1)+{r_{j}\over 2}\sum_{j=1}^{\infty}\left(\nu(a_{j},r_{j})-\delta_{a_{j}}\right)$
is a distribution of order 00 and locally of finite mass, where $\nu(a_{j},r_{j})$ is the Haar measure on the circle $S(a_{j},r_{j})$. The smoothing process was done in [2] as follows: as in the Appendix in [2], let
@@ 3. The counting function and the $k$-truncated counting function are defined by
$H(r):=4\pi\int_{0}^{r}h(s-c){ds\over s},$
where the constant $c\in(0,1/2)$ is chosen such that $H$ agrees with $\log^{+}r$ outside $[{1\over 2},{3\over 2}]$. Note that $H$ is an absolute function independent of $a_{j}$. This function $H(r)$ is smooth of $\log^{+}r$ with $0\leq H(r)-\log^{+}r\leq\log^{+}{3\over 2}<{1\over 2}$ and $H(r)=\log^{+}r$ when $H(r)\geq\log{3\over 2}$. Define
$\tau:=\log\sigma:=H(z)+\sum_{j=1}^{\infty}r_{j}H\left({r_{j}\over|z-a_{j}|}\right).$
Then $\sigma\geq\hat{\sigma}$ and the difference $\sigma-\hat{\sigma}$ is supported on
$\text{Supp}(\sigma-\hat{\sigma})\subset U:=\left(A\left(0,{1\over 2},{3\over 2}\right)\backslash{\mathcal{E}}\right)\bigcup_{j=1}^{\infty}A\left(a_{j},{1\over 2}r_{j},{3\over 2}r_{j}\right),$
where $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j}):=\{z\in{\mathcal{Y}}:{1\over 2}r_{j}\leq|z-a_{j}|\leq{3\over 2}r_{j}\}$ are pairwise disjoint annuli. For $z\not\in\cup_{j=1}^{\infty}\overline{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}$, $\log^{+}\left|{r_{j}\over z-a_{j}}\right|=H\left(\left|{r_{j}\over z-a_{j}}\right|\right)=0$ for each $j$. Thus $\tau(z)=H(|z|)$ and $\hat{\tau}(z)=\log^{+}|z|$, which are eqaual when $\tau(z)\geq\log{3\over 2}$. Let $B_{r}^{\sigma}:=\{z\in{\mathcal{Y}}~{}|~{}\sigma(z)<r\}$ be the $\sigma$-ball of radius $r$. Then the above discussion implies that
(3.2) $B_{t}^{\sigma}\subset{\mathbb{D}}_{t}~{}~{}~{}~{}(t\geq 1)~{}~{}~{}\mbox{and}~{}~{}~{}~{}~{}{\mathbb{D}}_{t}\backslash\left(\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})\right)\subset B_{t}^{\sigma}~{}~{}~{}~{}(t\geq{3\over 2}).$
We now show that the smooth 2-form $dd^{c}\sigma$ defines an order 0 distribution of finite mass on ${\mathcal{Y}}$ (see Lemma 3.2 in [2]) so that Parabolic Logarithmic Derivative Lemma (Lemma 9) holds for $\sigma$. We use the formula that, in polar coordinates $z=a+re^{i\theta}$ for some $a\in{\mathbb{C}}$, for a smooth function $\phi$, one has
@@
The smooth 2-form $dd^{c}\sigma$ is supported on $U$ since $\log\sigma$ is harmonic elsewhere. On the annulus $A(0,{1\over 2},{3\over 2})$, in polar coordinates $z=re^{i\theta}$ one has (see the Appendix in [2])
$dd^{c}H(|z|)=d\left({r\over 4\pi}{\partial H(r)\over\partial r}d\theta\right)=d(h(r-c)d\theta)=O(1)dr\wedge d\theta=O(1)rdr\wedge d\theta$
since ${1\over 2}\leq r\leq{3\over 2}$ is bounded. Thus $dd^{c}H(|z|)$ is of finite mass on $A(0,{1\over 2},{3\over 2})\backslash{\mathcal{E}}$. On the annulus $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})$,
$dd^{c}\left(r_{j}H\left({r_{j}\over|z-a_{j}|}\right)\right)$ $=d\left({r\over 4\pi}{\partial r_{j}H\left({r_{j}\over r}\right)\over\partial r}d\theta\right)$
$=d\left(-r_{j}h\left({r_{j}\over r}-c\right)d\theta\right)=O(r_{j})dr\wedge d\theta=O(1)rdr\wedge d\theta.$
@@
Since $\sum_{j=1}^{\infty}r^{2}_{j}\leq\sum_{j=1}^{\infty}r_{j}<+\infty$, the support $U$ is of finite Lebesgue measure. We conclude that $dd^{c}\log\sigma$ is of finite mass.
Using the above computations with ${\mathcal{E}}:=f^{-1}(D)$, we recall the following important result due to Z. Chen, D.T. Huynh, R. Sun and S.Y. Xie [2].

### v2

@@ 2.1. The Rational Case
(2.8) $N(\widetilde{G}(u_{1},\ldots,u_{n}))-N^{(1)}(\widetilde{G}(u_{1},\ldots,u_{n}))\leq 2\epsilon\deg\mathbf{f}.$
@@ 2.2. Motivation for the proof of the Main Theorem
In Section 2.1, the key to proving Theorem 6 is the use of the set $S$ (depending on the map) that contains all zeros and poles of each coordinate function of the map. The crucial observation is that all coordinate functions then become $S$-units, allowing us to apply Theorem 5. Importantly, the constants and the exceptional set obtained in Theorem 5 are independent of $S$, which makes the argument work uniformly for all rational maps.
To extend this viewpoint to the analytic setting, we let $\mathcal{E}_{f}$ denote the set of all zeros of $f_{i}$, $0\leq i\leq n$, and define
$\mathcal{Y}_{f}:=\mathbb{C}\setminus\mathcal{E}_{f},$
where $f=(f_{0},\ldots,f_{n})$ is a holomorphic map such that the entire functions $f_{i}$, $0\leq i\leq n$, have no common zeros. Then $\mathcal{Y}_{f}$ is a parabolic open Riemann surface equipped with a smooth exhaustion function $\sigma$ for which the smooth $(1,1)$-form $dd^{c}\!\log\sigma$ has finite total mass on $\mathcal{Y}_{f}$. Although $\sigma$ may not be a parabolic exhaustion for ${\mathcal{Y}}_{f}$, Chen et al. (see [2]) observed that, after appropriate modifications, the key ingredients of Nevanlinna theory still admit analogues in this setting.
@@ 3. Nevanlinna theory and the GCD theorem on parabolic Riemann surfaces
3.1. Parabolic Riemann surfaces and Jensen’s formula
@@
Let us first briefly review the construction of $\sigma$ given in Lemma 3.1 and 3.2 of [2]. Let $\{\mathcal{Y}$$\mathbb{C}$$\backslash\mathcal{E}$, where ${\mathcal{E}}:=\{a_{j}\}_{j=1}^{\infty}$ is a discrete countable set of points in $\mathbb{C}$}. Up to translation and dilation on $\mathbb{C}$}, we may assume that $|a_{j}|>2$ for all $j\geq 1$. We also assume that the radii satisfy the following conditions: (i) ${\mathbb{D}}(a_{j},2r_{j})$ for all $j$ and the disc ${\mathbb{D}}_{2}$ are disjoint, and (ii) the sum $\sum_{j\geq 1}r_{j}<+\infty$.
The exhaustion function $\sigma$ on $\mathcal{Y}$} is obtained by smoothing the piecewise smooth exhaustion function $\hat{\sigma}$ given by
(3.1) $\log\hat{\sigma}:=\log^{+}|z|+\sum_{j\geq 1}r_{j}\log^{+}{r_{j}\over|z-a_{j}|}.$
Note that $\hat{\tau}:=\log\hat{\sigma}$ takes values in $0,+\infty)$, is continuous on $\mathcal{Y}$} and is smooth (indeed harmonic) outside the circle $S(0,1):=\{|z|=1\}$ and the disjoint circles $S(a_{j},r_{j}):=\{|z-a_{j}|=r_{j}\}$. By the Poincaré-Lelong formula (see, e.g., [17]), it is clear that
$dd^{c}\hat{\tau}={1\over 2}\nu(0,1)+{1\over 2}\sum_{j=1}^{\infty}r_{j}\left(\nu(a_{j},r_{j})-\delta_{a_{j}}\right)$
is a distribution of order 0 and locally of finite mass, where $\nu(a_{j},r_{j})$ is the Haar measure on the circle $S(a_{j},r_{j})$.
The smoothing process goes as follows (see [2, Appendix]): Let
@@
(3.2) $H(r):=4\pi\int_{0}^{r}h(s-c){ds\over s},$
where the constant $c\in(0,1/2)$ is chosen such that $H$ agrees with $\log^{+}r$ outside ${1\over 2},{3\over 2}]$. Note that $H$ is an absolute function independent of $a_{j}$. This function $H(r)$ is smooth of $\log^{+}r$ with $0\leq H(r)-\log^{+}r\leq\log^{+}{3\over 2}<{1\over 2}$ and $H(r)=\log^{+}r$ when $H(r)\geq\log{3\over 2}$. Define
(3.3) $\tau:=H(|z|)+\sum_{j=1}^{\infty}r_{j}H\left({r_{j}\over|z-a_{j}|}\right),~~~~\sigma=\exp(\tau).$
Then, as stated in [2, Lemma 3.1], $\sigma\geq\hat{\sigma}$ and the difference $\sigma-\hat{\sigma}$ is supported on
(3.4) $\text{Supp}(\sigma-\hat{\sigma})\subset U:=\left(A\left(0,{1\over 2},{3\over 2}\right)\backslash{\mathcal{E}}\right)\bigcup\left(\bigcup_{j=1}^{\infty}A\left(a_{j},{1\over 2}r_{j},{3\over 2}r_{j}\right)\right),$
where $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j}):=\{z\in{\mathcal{Y}}:{1\over 2}r_{j}\leq|z-a_{j}|\leq{3\over 2}r_{j}\}$ are pairwise disjoint annuli. For $z\not\in\cup_{j=1}^{\infty}\overline{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}$, $\log^{+}\left|{r_{j}\over z-a_{j}}\right|=H\left(\left|{r_{j}\over z-a_{j}}\right|\right)=0$ for each $j$. Thus $\tau(z)=H(|z|)$ and $\hat{\tau}(z)=\log^{+}|z|$, which are equal when $\tau(z)\geq\log{3\over 2}$. Hence $\sigma=\hat{\sigma}$ when $\sigma\geq{3\over 2}$. We say that such $\sigma$ is associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$ by assuming that $\{r_{j}\}$ satisfy (i) and (ii) above. Note that $\sigma$ is not necessarily a parabolic exhaustion function because $dd^{c}\log\sigma$ may not be compactly supported.
Denote by $B^{\sigma}_{r}:=\{y\in{\mathcal{Y}}:\sigma(y)<r\}$ and by $S^{\sigma}_{r}:=\{y\in{\mathcal{Y}}:\sigma(y)=r\}$. Denote by $\mu_{r}$ the measure induced by the differential $d^{c}\log\sigma|_{S^{\sigma}_{r}}$ and write $d\mu_{r}=d^{c}\log\sigma|_{S^{\sigma}_{r}}$. Note that, from the above discussion (see also [2]),
(3.5) $B_{t}^{\sigma}\subset{\mathbb{D}}_{t}~~~~\mbox{for}~t\geq 1,~~~\mbox{and}~~~~~{\mathbb{D}}_{t}\backslash\left(\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})\right)\subset B_{t}^{\sigma}~~~~\mbox{for}~t\geq{3\over 2}.$
Lemma 8 (Lemma 3.2 in [2]).
The smooth 2-form $dd^{c}\log\sigma$ defines an order 0 distribution of finite mass on $\mathcal{Y}$}.
Proof.
We use the formula: in polar coordinates $z=a+re^{i\theta}$, for some $a\in{\mathbb{C}}$, and for a smooth function $\phi$, one has
@@
The smooth 2-form $dd^{c}\log\sigma$ is supported on $U$, where $U$ is given in (3.4). On the annulus $A(0,{1\over 2},{3\over 2})$, in polar coordinate $z=re^{i\theta}$, by (3.2) one has (see [2, Appendix])
(3.6) $dd^{c}\tau=dd^{c}H(|z|)=d\left({r\over 4\pi}{\partial H(r)\over\partial r}d\theta\right)=d(h(r-c)d\theta)=h^{\prime}\left(r-c\right)dr\wedge d\theta=O(1)dr\wedge d\theta.$
Thus $dd^{c}\tau$ has finite mass on $A(0,{1\over 2},{3\over 2})$. On the annulus $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})$, since ${\mathbb{D}}(a_{j},2r_{j})$ for all $j$ and the disc ${\mathbb{D}}_{2}$ are disjoint, we have similarly that $\tau=r_{j}H\left({r_{j}\over|z-a_{j}|}\right)$ on $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})$. Therefore,
$dd^{c}\tau$ $=dd^{c}\left(r_{j}H\left({r_{j}\over|z-a_{j}|}\right)\right)$
(3.7) $=d\left({r\over 4\pi}{\partial r_{j}H\left({r_{j}\over r}\right)\over\partial r}d\theta\right)=\left({r_{j}\over r}\right)^{2}h^{\prime}\left({r_{j}\over r}-c\right)dr\wedge d\theta=O(1)dr\wedge d\theta.$
@@
Since $\sum_{j=1}^{\infty}r_{j}<+\infty$, the support $U$ is of finite Lebesgue measure, this proves the lemma. ∎
We now derive the parabolic Jensen’s formula in our setting. For any function $\upsilon:{\mathcal{Y}}\rightarrow(-\infty,+\infty)$ such that $dd^{c}\upsilon$ is of order 0, the parabolic Jensen’s formula (see [16, Remark 3.2, p. 22]) is stated as
(3.8) $\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}dd^{c}\upsilon=\int_{S_{r}^{\sigma}}\upsilon d\mu_{r}-\int_{B_{r}^{\sigma}}\upsilon dd^{c}\tau\quad(r>1).$
Let $g$ be meromorphic on $\mathbb{C}$}. In (3.8), by letting $\upsilon:=\log|g|$, and also using the Poincaré-Lelong formula, we get the following lemma.
Lemma 9 (Parabolic Jensen’s formula).
Let $g$ be a meromorphic function on $\mathbb{C}$}. Then
$\int_{S_{r}^{\sigma}}\log|g|d\mu_{r}=N_{g,\sigma}(0,r)-N_{g,\sigma}(\infty,r)+c_{g,\sigma}(r),$
where $N_{g,\sigma}(0,r):=\int_{1}^{r}\sum_{z\in B_{t}^{\sigma}}\max\{0,{\rm ord}_{z}g\}{dt\over t},$ $N_{g,\sigma}(\infty,r):=N_{\frac{1}{g},\sigma}(0,r)$, and
(3.9) $c_{g,\sigma}(r):=\int_{B_{r}^{\sigma}}\log|g|dd^{c}\tau.$
Denote by
(3.10) $c^{+}_{g,\sigma}(r):=\int_{B_{r}^{\sigma}}\log^{+}|g|dd^{c}\tau.$
Lemma 10.
Let $g$ be a meromorphic function on $\mathbb{C}$}. Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$. Then we can choose $r_{j}$ small enough such that there exists a positive constant $C_{g}$ depending on $g$ with $c^{+}_{g,\sigma}(r)\leq C_{g}$ for any $r>0$.
Proof.
By using (3.5) for $t\geq 2$, and notice that $dd^{c}\tau$ is supported on $U$ where $U$ is given in (3.4), we have
[…]
(3.12) $=\int_{A(0,{1\over 2},{3\over 2})}\log^{+}|g|dd^{c}\tau+\int_{\cup_{j\geq 1}A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})\cap{\mathbb{D}}_{t}}\log^{+}|g|dd^{c}\tau.$
Notice that $\log^{+}|g|$ is locally integrable around the zeros or poles of $g$ and by (3.6), we get
$\int_{A(0,{1\over 2},{3\over 2})}\log^{+}|g|dd^{c}\tau<M_{1},$
where $M_{1}$ is a positive constant depending on $g$.
We now estimate the integration over $A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})$. If $g(a_{j})\not$$\infty$. Then $\log^{+}|g|\leq C_{1}$ for some $C_{1}>0$ on ${\mathbb{D}}(a_{j},{3\over 2}r_{j})$. Then, by (3.1),
$\int_{A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})}\log^{+}|g|dd^{c}\tau\leq C_{1}\cdot\frac{9\pi}{2}C_{h}r_{j}<{1\over 2^{j}},$
by further shrinking $r_{j}$, where $C_{h}:=\max_{\frac{1}{2}\leq s\leq\frac{3}{2}}|h^{\prime}(s-c)|.$ If $g(a_{j})=\infty$. Then, locally, $g(z)=(z-a_{j})^{k_{j}}g_{j}(z)$ with $g_{j}(a_{j})\not=0$. Thus $\log^{+}|g|\leq C_{2}\log^{+}{1\over|z-a_{j}|}$ for some $C_{2}>0$ on ${\mathbb{D}}(a_{j},{3\over 2}r_{j})$. We have similarly
$\int_{A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})}\log^{+}|g|dd^{c}\tau\leq\frac{9\pi}{2}C_{2}C_{h}\int_{{1\over 2}r_{j}}^{{3\over 2}r_{j}}\log rdr\leq\frac{9\pi}{2}C_{2}C_{h}r_{j}^{\delta},$
for some $1<\delta<2$. Again, we can choose $r_{j}$ sufficiently small such that
$\int_{A(a_{j},{1\over 2}r_{j},{3\over 2}r_{j})}\log^{+}|g|dd^{c}\tau<{1\over 2^{j}}.$
Therefore, with this choice of the radii $r_{j}$, we obtain from (3.11) that $c^{+}_{g,\sigma}(r)\leq M_{1}+\sum_{j=1}^{\infty}{1\over 2^{j}}=M_{1}+1$. ∎
Note that the shrinking process of $r_{j}$ in the above lemma depends on the given function $g$. So we must ensure in our arguments that such shrinking occurs only finitely many times. Throughout the rest of the paper, we will keep shrinking the radii $r_{j}$ whenever necessary. The $\sigma$ should be understood as being adjusted accordingly, although we will continue to denote it by the same symbol.
3.2. The characteristic functions in the parabolic setting
We continue to let $\{\mathcal{Y}$$\mathbb{C}$$\backslash\mathcal{E}$ with $\sigma$ being the exhaustion associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$, constructed in Section 3.1. Let $f:{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic map. We briefly recall some standard notations in Nevanlinna theory in the parabolic context (for details, see [2], [16]).
Let $D=\{Q=0\}$ be a hypersurface in $\mathbb{P}^{n}({\mathbb{C}})$, where $Q\in\mathbb{C}[x_{0},\ldots,x_{n}]$ is a homogeneous polynomial of degree $d$.
  1. 1.
    The proximity function of $f$ with respect to $D$ is defined as
    $m_{f,\sigma}(D,r)=\int_{S_{r}^{\sigma}}\lambda_{D}(f)d\mu_{r},$
    $\lambda_{D}({\bf x})=\log{\|{\bf x}\|^{d}\over|Q({\bf x})|},~~{\bf x}=[x_{0}:\cdots:x_{n}]\in\mathbb{P}^{n}({\mathbb{C}}),~~~\|{\bf x}\|:=\max\{|x_{0}|,\dots,|x_{n}|\}.$
  2. 2.
    The Nevanlinna characteristic (height) function is defined by
    $T_{f,\sigma}(r):=\int_{S_{r}^{\sigma}}\log\|{\bf f}\|d\mu_{r},~~~~r>1,$
    where ${\bf f}=(f_{0},\ldots,f_{n})$ is a reduced representation of $f$. The Ahlfors-Shimizu characteristic (height) function is defined by
    $T^{A}_{f,\sigma}(r)=\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}f^{*}\omega_{FS},~~~~r>1,$
    where $\omega_{FS}$ denotes the Fubini-Study form. Note that $T_{f,\sigma}(r)$ depends on the choices of the representation $\bf f$} of $f$, while $T^{A}_{f,\sigma}(r)$ does not depend on the representation. Therefore, when we write $T_{f,\sigma}(r)$, we always assume that we have fixed a reduced a representation ${\bf f}=(f_{0},\dots,f_{n})$ of $f$.
  3. 3.
    The counting function and the $k$-truncated counting function are defined by
    $N_{f,\sigma}(D,r):=\int_{1}^{r}\sum_{z\in B_{t}^{\sigma}}{\rm ord}_{z}f^{*}D{dt\over t},~~~~~~~\mbox{and}~~N_{f,\sigma}^{(k)}(D,r):=\int_{1}^{r}\sum_{z\in B_{t}^{\sigma}}\mathop{\mathrm{min}}\{k,{\rm ord}_{z}f^{*}D\}{dt\over t}.$
  4. 4.
    The counting function extends naturally to the sub-schemes, as well as the notion $N_{\text{gcd},\sigma}(f,g)(r)$, where $f,g$ are meromorphic functions on $\mathcal{Y}$. See [13].
We also use the standard notation $T_{f}(r),m_{f}(D,r),n_{f}(D,t)$ and $N_{f}(D,r)$ of the Nevanlinna functions associated to a holomorphic map $f:{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$. In this case, the Ahlfors-Shimizu and the classical Nevanlinna characteristic function agree up to an additive constant (see [14, Theorem 1.1.19]), i.e., $T_{f}(r)-T^{A}_{f}(r)=O(1).$
Theorem 11 (The parabolic First Main Theorem).
Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be as above. Let $D=\{Q=0\}$ be a hypersurface in $\mathbb{P}^{n}({\mathbb{C}})$, where $Q\in\mathbb{C}[x_{0},\ldots,x_{n}]$ is a homogeneous polynomial of degree $d$. Let $f=(f_{0},\ldots,f_{n}):{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic curve, where $f_{0},\ldots,f_{n}$ are entire functions without common zero. Suppose that ${f}({\mathcal{Y}})\not\subset\mbox{Supp}(D)$. Then
$dT_{f,\sigma}(r)=m_{f,\sigma}(D,r)+N_{f,\sigma}(D,r)+c_{Q({\bf f}),\sigma}(r).$
Proof.
Following the definition of Nevanlinna characteristic function, we have
$dT_{f,\sigma}(r)=\int_{S_{r}^{\sigma}}\log\|{\bf f}\|^{d}d\mu_{r}=\int_{S_{r}^{\sigma}}\log{\|{\bf f}\|^{d}\over|Q({\bf f})|}d\mu_{r}+\int_{S_{r}^{\sigma}}\log{|Q({\bf f})|}d\mu_{r}.$
Hence the assertion follows from Lemma 9 together with the definition of the proximity function. ∎
In the case of a meromorphic function $g$ on $\mathcal{Y}$}, write $g=g_{0}/g_{1}$, where $g_{0}$ and $g_{1}$ are entire functions without common zeros. We regard the function $g$ as the map $g:=[g_{0}:g_{1}]:{\mathcal{Y}}\rightarrow{\mathbb{P}}^{1}({\mathbb{C}})$. By taking $Q(z_{0},z_{1})=z_{1}$ (so $D=\{z_{1}=0\}=\infty$), we get
$m_{g,\sigma}(r):=m_{g,\sigma}(\infty,r)=\int_{S_{r}^{\sigma}}\log{\max\{|g_{0}|,|g_{1}|\}\over|g_{1}|}d\mu_{r}=\int_{S_{r}^{\sigma}}\log^{+}|g|d\mu_{r}+O(1).$
The above $m_{g,\sigma}(r)$ is well-defined, i.e., it is independent of the choice of $g_{0}$ and $g_{1}$. However, the Nevanlinna height function $T_{g,\sigma}(r)=\int_{S_{r}^{\sigma}}\log\max\{|g_{0}|,|g_{1}|\}d\mu_{r}$ is not well-defined. Instead we introduce a new height function that is well-defined:
(3.13) $T^{M}_{g,\sigma}(r)=m_{g,\sigma}(\infty,r)+N_{g,\sigma}(\infty,r).$
From the parabolic Jenen’s formula above,
(3.14) $T^{M}_{1/g,\sigma}(r)=T^{M}_{g,\sigma}(r)+c_{1/g,\sigma}(r).$
We now compare $T^{M}_{g,\sigma}(r)$ with $T^{A}_{g,\sigma}(r)$. From parabolic Jensen’s formula,
$T^{A}_{g,\sigma}(r)$ $=$ $\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}dd^{c}\log\max\{|g_{0}|,|g_{1}|\}=\int_{S_{r}^{\sigma}}\log\max\{|g_{0}|,|g_{1}|\}d\mu_{r}$
$-\int_{B_{r}^{\sigma}}\log\max\{|g_{0}|,|g_{1}|\}dd^{c}\tau$
$=$ $\int_{S_{r}^{\sigma}}\log^{+}|g|d\mu_{r}+\int_{S_{r}^{\sigma}}\log|g_{1}|d\mu_{r}-\int_{B_{r}^{\sigma}}\log\max\{|g_{0}|,|g_{1}|\}dd^{c}\tau$
$=$ $T^{M}_{g,\sigma}(r)+{1\over 2}\int_{B_{r}^{\sigma}}\log\left(1/(1+|g|^{2})\right)dd^{c}\tau=T^{M}_{g,\sigma}(r)+{1\over 2}c_{1/(1+|g|^{2}),\sigma}(r).$
Similarly to the proof of Lemma 10, by choosing $r_{j}$ small enough, there exists a constant $C>0$ with $c^{+}_{1/(1+|g|^{2}),\sigma}(r)\leq C$. Hence
(3.15) $|T^{M}_{g,\sigma}(r)-T^{A}_{g,\sigma}(r)|\leq C.$
It is important to note that these radii $r_{j}$ depend on the function $g$.